OpenStudy (anonymous):

Solve. (1/36)= 6^(x-3) A) x = -5 B) x = - 3/2 C) x = 7/2 D) x = 1

OpenStudy (anonymous):

Hint: ln both sides first in order to bring down (x-3)

OpenStudy (anonymous):

ln= natural log

OpenStudy (anonymous):

I'm not sure if bringing logs into this is entirely necessary. \[1/36 = 6^{-2}\] If you subtract that from both sides you get \[0=6^{x-3}-6^{-2}\] There's gotta be some way to solve for x from there, I just can't think of it right now.

OpenStudy (anonymous):

You have to use natural log. That or you just plug in values until you find the right one.

OpenStudy (anonymous):

try \(\frac{1}{36}=6^{-2}\)

OpenStudy (anonymous):

no logs involved \[6^{-2}=6^{x-3}\iff -2=x-3\]

OpenStudy (anonymous):

solve for \(x\) in one step by adding 3 to both sides of \[-2=x-3\] to get \[x=1\]

OpenStudy (anonymous):

Aha, that's right. I remember now =) If the bases are the same, you can set the exponents equal to each other.

OpenStudy (anonymous):

I think it's better to get people use to solving natural logs. It's only going to get more complicated further down the road.